Principles of Corporate Finance (12th Edition) , vhapter 25 task 8p: why do they call it pre-tax, then remove 35% tax from that? Like pre-tax is 400, then they take 400*0,65= 260
Answer #1

Pre tax means an income before the consideration of tax.

Once tax is considered then the income will be post tax.

Post tax= Pre tax income*(1-tax)

Pre tax income= $400

Tax rate= 35%

Post tax income= 400*(1-0.35)= 400*0.65= $260